Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
|
OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn |
Not logged in |
 |
|
Message ID: 144
Entry time: 19 Apr 2023, 23:20
|
Author: |
Laurent Mercadier, Martin Teichmann |
Group: |
SCS |
Subject: |
shift summary day 2 UP3463 |
|
|
Achievements
- found larger beamstop
- found center of pinhole and bender settings for beam size
- checked jitter of beam on grating, looks like jittering is large, to be further investigated offline
- checked that patterns are working for the night, to be continued by night crew
Problems
- Major problems with scanner Y, as during the previous runs. Reset axis, Reset and adjust the target position to actual position + move to get out of error.
- Problem with bunch pattern (see separate entry in Operations elog)
-
|